OpenStudy (anonymous):

Suppose Q and R are independent events. Find P(Q and R) P(Q) = 4/5, P(R) = 4/11

OpenStudy (anonymous):

Suppose Q and R are independent events. Find P(Q and R) P(Q) = 4/5, P(R) = 4/11